# Can NeRFs See without Cameras?

### v2

@@ 1 Introduction
NeRFs mildenhall2021nerf ; nerfInvProb ; mipnerf ; pixelnerf have delivered impressive results in solving inverse problems, resulting in 3D scene rendering. While NeRFs have mostly used pictures (from cameras or LIDARs) to infer a 3D scene, we ask if the core ideas can generalize to the case of wireless signals (such as RF or audio). Unlike camera pixels that receive line-of-sight (LoS) rays, a wireless receiver (e.g., a WiFi antenna on a smartphone) would receive a mixture of LoS and many reflections, called multipath. If the receiver moves, it receives a sequence of $N$ measurements. Using these $N$ wireless measurements, is it possible to learn a representation of the scene, such as the floorplan of the user’s home? An optical NeRF understandably fails since it is not equipped to handle multipath. This paper is focused on redesigning NeRFs so they can learn to image the environment, thereby solving the inverse problem from ambient wireless signals.
A growing body of work specnerf ; winert ; zhao2023nerf2 ; newrf ; luo2022learning ; AV-NeRF is investigating connections between NeRFs and wireless. While none have concentrated on imaging, NeRF2 zhao2023nerf2 and NeWRF newrf have augmented NeRFs to correctly synthesize WiFi signals at different locations inside an indoor space. However, correct synthesis is possible without necessarily learning the correct signal propagation models. We find that NeRFs with adequate model complexity can overfit a function to correctly predict signals at test locations, but this function does not embed the true behavior of multipath signal propagation. We re-design the NeRF’s objective function so that it learns the environment through line-of-sight (LoS) paths and reflections. This teaches the NeRF an implicit representation of the scene, which can then be utilized for various forward tasks, including WiFi signal prediction and ray tracing.
@@
To evaluate EchoNeRF, we train on $2.4$ GHz WiFi signals from NVIDIA’s Sionna simulator hoydis2023sionna , with floorplans from the Zillow’s Indoor Dataset (ZIND) cruz2021zillow . Results show consistent improvement over baselines in terms of the estimated floorplan’s IoU and F1 score. Qualitative results show visually legible floorplans without any post-processing. Applying forward functions on the floorplan, EchoNeRF can predict the received signal power for new $\langle Tx,Rx\rangle$ locations (outperforming existing baselines). Lastly, basic ray tracing explains the predictions, offering interpretability to its results.
@@ 2 Related Work and Research Scope
Wireless (WiFi) channel prediction using NeRFs. NeWRFnewrf and NeRF2zhao2023nerf2 are recent papers that have used NeRFs to predict the wireless channel impulse response (CIR) Tse_Viswanath_2005 at unknown locations inside a room. Drawing a parallel to optical NeRFs, a voxel’s color in optics becomes a voxel’s transmit power in wireless. The voxel’s density in optics remains the same in wireless, modeling how that voxel attenuates signals passing through it. NeRF2 and NeWRF assign transmit power and attenuation to each voxel such that they best explain the measured CIR. The authors explain that voxels assigned non-zero transmit power will be called virtual transmitters; these voxels represent the reflection points on the walls. However, many assignments are possible that fit the CIR training data, especially when the data is sparse. Fig. 2(a,b) illustrates 2 possible assignments. While the predicted CIRs could achieve low error for all such assignments, only one of the assignments will model the true reflections, forcing the virtual transmitters to be located on the wall surfaces (Fig. 2(b)). We have plotted NeRF2 and NeWRF’s assignment of voxel densities (see Fig. 2(c,d)) to confirm that the high accuracy in CIR prediction is not an outcome of correctly learning the wall layout. Our goal is to repair this important issue, i.e., assign voxel densities that obey the basic physics of wall reflections. Correct voxel assignment leads to the correct layouts, which then makes the (forward) CIR prediction easy.
@@
Neural radiance fields for audio. Another active line of research focuses on predicting room impulse response (RIR) for audio luo2022learning ; INRAS ; AV-NeRF ; brunetto2024neraf . Neural Acoustic Field (NAF) luo2022learning extended the classical NeRF to train on RIR measurements in a room and predict the RIR (magnitude and phase) at new $\langle Tx,Rx\rangle$ locations. NAF identified the possibility of overfitting to the RIR and proposed to learn, jointly, the local geometric features of the environment (as spatial latents) and the NAF parameters. The spatial latents embed floorplan information but a decoder needs to be trained using the partial floorplan data. EchoNeRF requires no floorplan supervision, and secondly, relies entirely on signal power (less informative than RIR) to estimate the floorplan.
@@
Follow up work are embracing more information about the surroundings (pictures NACF ; INRAS ; AV-NeRF ; fewshotvisual , LIDAR scans oufqir2020arkit , meshes and optical NeRFs brunetto2024neraf ), to boost RIR accuracy. Results are steadily improving, however, this sequence of ideas is unaligned with solving the core inverse problem. Our goal is to first invert the signal power to a floorplan, which can then enable CIR/RIR predictions.
@@
Modeling reflections in optical NeRFs Optical NeRFs have tackled reflections refneus ; nerfds ; flipnerf for synthesizing glossy surfaces and mirrors, and for re-lighting rudnev2022nerf ; relight . NeRFRen nerfren proposes to decompose a viewed image into a transmitted and a reflected component. Ref-NeRF verbin2022refnerf also focuses on reflections through a similar decomposition of the transmitted and reflected color, however, the reflected color is modeled as a function of the viewing angle and the surface-normal, resulting in accurate models of specular reflection. Several other papers have developed similar ideas boss2021nerd ; zhang2021nerfactor and the core insight centers around solving a two-component decomposition problem. EchoNeRF faces the challenge of not knowing the number of rays adding up from all possible directions in the environment. Hence, EchoNeRF must solve a many-component decomposition problem by leveraging the physics of multipath signal propagation.
@@ 3 EchoNeRF Model
where $\psi_{LoS}$ is the power from the direct line-of-sight (LoS) path, and $\psi_{ref_{k}}$ is the aggregate power from all $k^{th}$ order reflections (i.e., all signal paths that underwent exactly $k$ reflections before arriving at the $Rx$)  [^1] This model is an approximation since it relies on averaging measurements as explained in Appendix B. However, for this sensing application involving moving receivers, and when using wideband-multipath systems such as WiFi, this approximation holds well. We assume $M$ fixed transmitters and move the $Rx$ to $N$ known locations and measure $\psi$ at each of them. EchoNeRF accepts $M\times N$ measurements as input and outputs the 2D floor-plan $F$, a binary matrix of size $L\times L$, where $L$ denotes the maximum floorplan length.
@@ 3.1 The LoS Model
Friss’ equation antennaBook from electromagnetics models the free-space received power as $P_{r}=\frac{K}{d^{2}}$ where $d$ is the distance of signal propagation, and $K$ is a product of transmit-power, wavelength, and antenna-related constants antennaBook . We model this free-space (LoS) behavior in the NeRF framework through the following equation.
@@ 3.2 The Reflection Model
We address this by modeling $\omega_{j}$ as a discrete value, i.e., $\omega_{j}\in\{1,2\dots K_{\omega}\}$. Larger $K_{\omega}$ is needed when the environment has complex surface orientations; however, most floorplans exhibit perpendicular walls manhattan1 ; manhattan2 ; manhattan3 and $K_{\omega}=4$ is adequate. Once $\omega_{j}$ becomes discrete, the voxels that can produce plausible reflections become far fewer – we call this the “plausible set” $\mathcal{V}$. Fig.  4 visualizes $\mathcal{V}$ and shows $3$ out of many plausible reflections from voxels with $\omega_{j}=45^{\circ}$. Eq. 3 sums up the power from all reflections that occur on the plausible set:
@@ 3.4 Multi-stage Training
$\blacksquare$ Regularization: Floorplans demonstrate significant local similarity in orientation, hence we penalize differences in orientation among neighbors, using a regularization (Eq. 8) similar to Total Variation TV_reg . This can be achieved without additional computational cost to the neural network by directly utilizing voxel orientations obtained from each ray.
@@ 4 Experiments
Floor plan and Wireless Simulation Dataset. Floorplans are drawn from the Zillow Indoor Dataset  cruz2021zillow . In each floorplan, we use the $A^{*}$ algorithm  astar to generate a walking trajectory that traverses all rooms. We use the NVIDIA Sionna RT hoydis2022sionna ; hoydis2023sionna – a ray tracer for radio propagation modeling – to compute the ground truth signal power (also known as received signal strength index (RSSI)). We randomly place $M$ Txs, one in each room, denoted as $T_{m}$. To simulate omnidirectional transmissions at 2.42.4GHz from each Tx location, we shoot $10^{7}$ rays into the given floorplan. For receiver locations, we sample the user trajectory at a fixed time interval to obtain $N$ Rx locations, denoted as $R_{n}$. Sionna accounts for specular reflections and refraction when these rays interact with walls in the specified floor plan; we use the default materials for the walls. As with most WiFi simulators, Sionna does not model signal penetration through walls – this means that a Tx and Rx located on opposite sides of a wall will not receive any RSSI. Overall, we gather $M\times N$ RSSI measurements $(T_{m},R_{n},\psi_{m,n})$ that serve as input to EchoNeRF.
@@
    NeRF2 zhao2023nerf2 : Models WiFi reflections via virtual transmitters to predict channel impulse response (CIR).
@@
    Heatmap Segmentation liu2012review : Interpolates CIR across the whole floorplan and applies an image segmentation algorithm (on the interpolated RSSI heatmap) to isolate each room. Essentially, the algorithm identifies the contours of sharp RSSI change since such contours are likely to correspond to walls. Implementation details are included in the Appendix E.4.
@@
where $WP$ denotes the set of predicted wall pixels and $WP^{*}$ denotes the true wall pixels. This is a harsh metric given wall pixels are a small fraction of the total floorplan; if a predicted wall is even offset by one pixel from the true wall, the Wall_IoU drops significantly. IoU iou has often been defined in terms of room pixels (instead of wall pixels); this is an overestimate in our opinion, since predicting even an empty floorplan results in an impressively high IoU.
(B) F1 score metrics : Defined as $F1=\frac{2\times P\times R}{P+R}$, where $P$ is the precision and $R$ is the recall of the bitmap. $P$ and $R$ are defined based on wall pixels, similar as above.
@@ 4.1 Overall Summarized Results
Heatmap Seg. Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
NeRF2 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
EchoNeRF Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 5: Qualitative comparison of ground truth floorplans against baselines. The bottom two rows show floorplans learnt by EchoNeRF_LoS and EchoNeRF with sharper walls and boundaries.
@@ 5 Follow ups and Conclusion
Follow-ups. (1) The ability to model $2^{nd}$ order reflections will boost EchoNeRF’s accuracy, allowing it to sharpen the scene and decode smaller objects. For short range applications, such as non-intrusive medical imaging, $2^{nd}$ and $3^{rd}$ order reflections would be crucial. This remains an important direction for follow-on research. (2) Extending EchoNeRF to 3D floorplans is also of interest, and since it is undesirable to increase the number of measurements, effective 3D priors, or 2D-to-3D post-processing, may be necessary. Such post-processing tools exist cedreo2025 but we have not applied them since our goal is to improve NeRF’s inherent inverse solver. (3) Finally, EchoNeRF floorplans can offer valuable spatial context to Neural RIR synthesizers like zhao2023nerf2 ; newrf ; brunetto2024neraf ; AV-NeRF ; NACF . Synthesized RIR could in-turn aid EchoNeRF’s floorplan inference, forming the basis for an alternating optimization strategy. We leave these ideas to follow-up research.
Conclusion. In summary, we re-design the NeRF framework so it can learn to "see" its environment through line-of-sight (LoS) paths and reflections. This work takes a step towards neural wireless imaging with room for further research and varied downstream applications.

### v3

# Can NeRFs “See” without Cameras?

@@ 1 Introduction
NeRFs Mildenhall et al. (2021); Wang et al. (2023); Barron et al. (2021); Yu et al. (2021) have delivered impressive results in solving inverse problems, resulting in 3D scene rendering. While NeRFs have mostly used pictures (from cameras or LIDARs) to infer a 3D scene, we ask if the core ideas can generalize to the case of wireless signals (such as RF or audio). Unlike camera pixels that receive line-of-sight (LoS) rays, a wireless receiver (e.g., a WiFi antenna on a smartphone) would receive a mixture of LoS and many reflections, called multipath. If the receiver moves, it receives a sequence of $N$ measurements. Using these $N$ wireless measurements, is it possible to learn a representation of the scene, such as the floorplan of the user’s home? A vanilla NeRF understandably fails since it is not equipped to handle multipath. This paper is focused on redesigning NeRFs so they can learn to image the environment, thereby solving the inverse problem from ambient wireless signals.
A growing body of work Basu and Chakraborty (2024); Orekondy et al. (2023); Zhao et al. (2023); Lu et al. (2024); Luo et al. (2022); Liang et al. (2023a) is investigating connections between NeRFs and wireless. While none have concentrated on imaging, NeRF2 Zhao et al. (2023) and NeWRF Lu et al. (2024) have augmented NeRFs to correctly synthesize WiFi signals at different locations inside an indoor space. However, correct synthesis is possible without necessarily learning the correct signal propagation models. We find that NeRFs with adequate model complexity can overfit a function to correctly predict signals at test locations, but this function does not embed the true behavior of multipath signal propagation. We re-design the NeRF’s objective function so that it learns the environment through line-of-sight (LoS) paths and reflections. This teaches the NeRF an implicit representation of the scene, which can then be utilized for various forward tasks, including WiFi signal prediction and ray tracing.
@@
To evaluate EchoNeRF, we train on $2.4$ GHz WiFi signals from NVIDIA’s Sionna simulator Hoydis et al. (2023), with floorplans from the Zillow’s Indoor Dataset (ZIND) Cruz et al. (2021). Results show consistent improvement over baselines in terms of the estimated floorplan’s IoU and F1 score. Qualitative results show visually legible floorplans without any post-processing. Applying forward functions on the floorplan, EchoNeRF can predict the received signal power for new $\langle Tx,Rx\rangle$ locations (outperforming existing baselines). Lastly, basic ray tracing explains the predictions, offering interpretability to its results.
@@ 2 Related Work and Research Scope
Wireless (WiFi) channel prediction using NeRFs. NeWRFLu et al. (2024) and NeRF2Zhao et al. (2023) are recent papers that have used NeRFs to predict the wireless channel impulse response (CIR) Tse and Viswanath (2005) at unknown locations inside a room. Drawing a parallel to optical NeRFs, a voxel’s color in optics becomes a voxel’s transmit power in wireless. The voxel’s density in optics remains the same in wireless, modeling how that voxel attenuates signals passing through it. NeRF2 and NeWRF assign transmit power and attenuation to each voxel such that they best explain the measured CIR. The authors explain that voxels assigned non-zero transmit power will be called virtual transmitters; these voxels represent the reflection points on the walls. However, many assignments are possible that fit the CIR training data, especially when the data is sparse. Fig. 2(a,b) illustrates 2 possible assignments. While the predicted CIRs could achieve low error for all such assignments, only one of the assignments will model the true reflections, forcing the virtual transmitters to be located on the wall surfaces (Fig. 2(b)). We have plotted NeRF2 and NeWRF’s assignment of voxel densities (see Fig. 2(c,d)) to confirm that the high accuracy in CIR prediction is not an outcome of correctly learning the wall layout. Our goal is to repair this important issue, i.e., assign voxel densities that obey the basic physics of wall reflections. Correct voxel assignment leads to the correct layouts, which then makes the (forward) CIR prediction easy.
@@
Neural radiance fields for audio. Another active line of research focuses on predicting room impulse response (RIR) for audio Luo et al. (2022); Su et al. (2022); Liang et al. (2023a); Brunetto et al. (2024). Neural Acoustic Field (NAF) Luo et al. (2022) extended the classical NeRF to train on RIR measurements in a room and predict the RIR (magnitude and phase) at new $\langle Tx,Rx\rangle$ locations. NAF identified the possibility of overfitting to the RIR and proposed to learn, jointly, the local geometric features of the environment (as spatial latents) and the NAF parameters. The spatial latents embed floorplan information but a decoder needs to be trained using the partial floorplan data. EchoNeRF requires no floorplan supervision, and secondly, relies entirely on signal power (less informative than RIR) to estimate the floorplan.
@@
Follow up work are embracing more information about the surroundings (pictures Liang et al. (2023b); Su et al. (2022); Liang et al. (2023a); Majumder et al. (2022), LIDAR scans Oufqir et al. (2020), meshes and optical NeRFs Brunetto et al. (2024)), to boost RIR accuracy. Results are steadily improving, however, this sequence of ideas is unaligned with solving the core inverse problem. Our goal is to first invert the signal power to a floorplan, which can then enable CIR/RIR predictions.
@@
Modeling reflections in optical NeRFs Optical NeRFs have tackled reflections Ge et al. (2023); Yan et al. (2023); Seo et al. (2023) for synthesizing glossy surfaces and mirrors, and for re-lighting Rudnev et al. (2022); Toschi et al. (2023). NeRFRen Guo et al. (2022b) proposes to decompose a viewed image into a transmitted and a reflected component. Ref-NeRF Verbin et al. (2022) also focuses on reflections through a similar decomposition of the transmitted and reflected color, however, the reflected color is modeled as a function of the viewing angle and the surface-normal, resulting in accurate models of specular reflection. Recent proposals such as NeRF-Casting Verbin et al. (2024) and oRCA Tiwary et al. (2023) have further improved the models of multipath for glossy and mirrored surfaces, and others Boss et al. (2021); Zhang et al. (2021) have developed similar ideas, with the core insights centering around solving a two-component decomposition problem. EchoNeRF faces the challenge of not knowing the number of rays adding up from all possible directions in the environment. Hence, EchoNeRF must solve a many-component decomposition problem by leveraging the physics of multipath signal propagation.
In the context of LIDARs, PlatoNeRF Klinghoffer et al. (2024) and NeTF Shen et al. (2021) cope with unknown reflections, however, LIDARs have very high time resolution (due to high clock frequency), and is therefore able to assign the incoming rays to different time buckets. This temporal separation allows the NeRF to make separate measurements for different surfaces in the scene. Since EchoNeRF uses only signal power—a single scalar measurement that contains a mixture of the LoS and all the reflections—the inverse problem of recovering the voxel attributes from only the power is far more complex.
@@ 3 EchoNeRF Model
where $\psi_{LoS}$ is the power from the direct line-of-sight (LoS) path, and $\psi_{ref_{k}}$ is the aggregate power from all $k^{th}$ order reflections (i.e., all signal paths that underwent exactly $k$ reflections before arriving at the $Rx$)  [^1] This model is a simplification since it ignores the signal phase in estimating the received power. Appendix B shows that with a moving wide-band receiver, like WiFi, the approximation may be tolerable for a sensing application like EchoNeRF. We assume $M$ fixed transmitters and move the $Rx$ to $N$ known locations and measure $\psi$ at each of them. EchoNeRF accepts $M\times N$ measurements as input and outputs the 2D floor-plan $F$, a binary matrix of size $L\times L$, where $L$ denotes the maximum floorplan length.
@@ 3.1 The LoS Model
Friss’ equation Balanis (2016) from electromagnetics models the free-space received power as $P_{r}=\frac{K}{d^{2}}$ where $d$ is the distance of signal propagation, and $K$ is a product of transmit-power, wavelength, and antenna-related constants Balanis (2016). We model this free-space (LoS) behavior in the NeRF framework through the following equation.
@@ 3.2 The Reflection Model
We address this by modeling $\omega_{j}$ as a discrete value—multiples of $\frac{\pi}{K_{\omega}}$. Larger $K_{\omega}$ is needed when the environment has complex surface orientations; however, most floorplans exhibit perpendicular walls Zou et al. (2018); Coughlan and Yuille (1999); Guo et al. (2022a) and $K_{\omega}=4$ is adequate. Once $\omega_{j}$ becomes discrete, the voxels that can produce plausible reflections become far fewer – we call this the “plausible set” $\mathcal{V}$. Fig.  4 visualizes $\mathcal{V}$ and shows $3$ out of many plausible reflections from voxels with $\omega_{j}=135^{\circ}$. Eq. 3 sums up the power from all reflections that occur on the plausible set:
@@ 3.4 Multi-stage Training
$\blacksquare$ Regularization: Floorplans demonstrate significant local similarity in orientation, hence we penalize differences in orientation among neighbors, using a regularization (Eq. 8) similar to Total Variation Rudin et al. (1992). This can be achieved without additional computational cost to the neural network by directly utilizing voxel orientations obtained from each ray.
@@ 4 Experiments
Floorplan and Wireless Simulation Dataset. Floorplans are drawn from the Zillow Indoor Dataset  Cruz et al. (2021). In each floorplan, we use the $A^{*}$ algorithm  Hart et al. (1968) to generate a walking trajectory that traverses all rooms. We use the NVIDIA Sionna RT Hoydis et al. (2022, 2023) – a ray tracer for radio propagation modeling – to compute the ground truth signal power (also known as received signal strength index (RSSI)). We randomly place $M$ Txs, one in each room, denoted as $T_{m}$. To simulate omnidirectional transmissions at 2.42.4GHz from each Tx location, we shoot $10^{7}$ rays into the given floorplan. For receiver locations, we sample the user trajectory at a fixed time interval to obtain $N$ Rx locations, denoted as $R_{n}$. Sionna accounts for specular reflections and refraction when these rays interact with walls in the specified floor plan; we use the default materials for the walls. As with most WiFi simulators, Sionna does not model signal penetration through walls – this means that a Tx and Rx located on opposite sides of a wall will not receive any RSSI. Overall, we gather $M\times N$ RSSI measurements $(T_{m},R_{n},\psi_{m,n})$ that serve as input to EchoNeRF.
@@
    NeRF2 Zhao et al. (2023): Models WiFi reflections via virtual transmitters to predict channel impulse response (CIR).
@@
    Heatmap Segmentation Liu et al. (2012): Interpolates CIR across the whole floorplan and applies an image segmentation algorithm (on the interpolated RSSI heatmap) to isolate each room. Essentially, the algorithm identifies the contours of sharp RSSI change since such contours are likely to correspond to walls. Implementation details are included in the Appendix E.4.
@@
where $WP$ denotes the set of predicted wall pixels and $WP^{*}$ denotes the true wall pixels. This is a harsh metric given wall pixels are a small fraction of the total floorplan; if a predicted wall is even offset by one pixel from the true wall, the Wall_IoU drops significantly. IoU Rezatofighi et al. (2019) has often been defined in terms of room pixels (instead of wall pixels); this is an overestimate in our opinion, since predicting even an empty floorplan results in an impressively high IoU.
(B) F1 score Sokolova and Lapalme (2009): Defined as $F1=\frac{2\times P\times R}{P+R}$, where $P$ is the precision and $R$ is the recall of the bitmap. $P$ and $R$ are defined based on wall pixels, similar as above.
@@ 4.1 Overall Summarized Results
6cm. Heatmap Seg. Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
6cm. NeRF2 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
6cm. EchoNeRF Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 5: Qualitative comparison of ground truth floorplans against baselines. In the first row, red stars denote Tx locations and light gray dots denote Rx measurement locations. The bottom two rows show floorplans learnt by EchoNeRF_LoS (i.e., Stage 1) and EchoNeRF (i.e., Stage 2) with sharper walls and boundaries. More visualizations available at https://echonerf.github.io/
@@ 4.3 Relaxing Assumptions & Sensitivity Study
$\blacksquare$ Material Sensitivity. Table 4 shows the mean Wall_IoU of EchoNeRF on five different materials averaged across six scenes (shown in Fig. 5) Materials with higher reflectivity, such as concrete and glass, yield better performance than absorptive materials like wood. This is because more reflections allow better performance for EchoNeRFś reflection model.
$\blacksquare$ Robustness to RSSI error. We added Gaussian noise to the RSSI measurements with a mean equal to the noise floor (in dB) and a variance of $4$ dB. We vary the noise floor levels ranging from 80-80 dB to 130-130 dB across the 6 floorplans shown in Fig. 5. The SNR at a receiver is computed as the difference between the received signal power and the noise floor (e.g., a received power of 70-70 dB with a noise floor of 80-80 dB results in an SNR of $10$ dB). We report the mean Wall_IoU for EchoNeRF_LoS and EchoNeRF in Table 4. The performance drops with decreasing SNR; both EchoNeRF_LoS and EchoNeRF’s floorplans are still legible till $30$ dB, but below that (when noise power becomes comparable to the signal power) they break down, leading to missing and illegible walls. Our results are conservative; when we report a specific SNR level (e.g., 40dB), it represents the highest SNR (best-case) among all receivers in that scenario, meaning other receivers experience even lower SNRs. In practice, WiFi SNR ranges from 30-60dB depending on the distance from the router, with close-proximity measurements often exceeding 50-60dB. For average real-world SNR conditions around 45dB, the corresponding best-case SNR would be 60+dB, which aligns with the top rows where EchoNeRF demonstrates strong performance.
[…]
Concrete 0.251 0.371
Glass 0.236 0.364
Brick 0.232 0.357
Marble 0.226 0.328
Wood 0.227 0.311
Table 3: Sensitivity across materials.
[…]
$\infty$ (no noise) 0.251 0.371 Legible
60 0.246 0.336 Legible
50 0.231 0.292 Legible
40 0.226 0.298 Legible
30 0.207 0.241 Missing walls
10 0.090 0.140 Illegible
Table 4: Noise robustness across SNR levels.
@@ 5 Follow ups and Conclusion
Follow-ups. (1) The ability to model $2^{nd}$ order reflections will boost EchoNeRF’s accuracy, allowing it to sharpen the scene and decode smaller objects. For short range applications, such as non-intrusive medical imaging, $2^{nd}$ and $3^{rd}$ order reflections would be crucial. This remains an important direction for follow-on research. (2) Extending EchoNeRF to 3D floorplans is also of interest, and since it is undesirable to increase the number of measurements, effective 3D priors, or 2D-to-3D post-processing, may be necessary. Such post-processing tools exist Cedreo (2025) but we have not applied them since our goal is to improve NeRF’s inherent inverse solver. (3) Expanding evaluations beyond ZInD, which contains largely unfurnished, rectangular rooms, to richer datasets such as HM3D Ramakrishnan et al. (2021) and MVL Solarte et al. (2024) would also be valuable: furniture and clutter can introduce significant multipath effects that complicate RF signal modeling, and additional research is needed to understand these effects. We defer this to future work, as our focus in this paper is not so much to understand the limits of RF-based NeRFs, but to establish the feasibility of such frameworks. (4) Finally, EchoNeRF floorplans can offer valuable spatial context to Neural RIR synthesizers like Zhao et al. (2023); Lu et al. (2024); Brunetto et al. (2024); Liang et al. (2023a, b). Synthesized RIR could in-turn aid EchoNeRF’s floorplan inference, forming the basis for an alternating optimization strategy. We leave these ideas to follow-up research.
Conclusion. In summary, we re-design the NeRF framework so it can learn to "see" its environment by leveraging both line-of-sight (LoS) paths and multipath reflections. While such reflections bring to the sensor more information about the surroundings, their mixtures with the LoS path also complicates the core inverse problems. EchoNeRF takes a step towards solving this inverse problem, but also leaves room for further research in neural wireless imaging and varies downstream applications.
Acknowledgments
We would like to thank Prof. Shenlong Wang for his guidance in the initial phase of this work. We also thank the anonymous reviewers for their valuable feedback. This work was partially supported by NSF #2008338, #1909568, #2148583, and #MRI-2018966. This work used DELTA at NCSA through allocation CIS230230 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by U.S. National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296